40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 613     Entry time: Tue Jul 1 12:51:34 2008
Author: John 
Type: Summary 
Category: General 
Subject: IFO alignment 
Rana, Rob, Yoichi, John

The recent computer problems and MZ work had disturbed the alignment of the interferometer.

We adjusted the MC alignment back to nominal positions using old OSEM values. We then walked
the input beam to match the MC. Coupling into the interferometer has increased noticeably.
The rest of the IFO was then aligned to the new input beam.

Proceeding to full IFO locking we were able to engage the AO path and hand off CARM to SPOBDC.
Arm powers got up to 4.

If the new alignment proves successful we will centre all QPDs etc so we can easily return to
this state.
Attachment 1: align080701.png  106 kB  Uploaded Tue Jul 1 14:06:08 2008  | Show | Show all
ELOG V3.1.3-